OpenStudy (anonymous):

Dani solves the system of equations shown and finds that the solution is (-2,7) explain whether dani is correct. 7x-y=7 -3x+2y=8

OpenStudy (anonymous):

7x-y=7 -3x 2y=8 (7 x -2) -7 =7 -1-7 = 7 -8 =/= 7 (-3 x -2) + (2 x 7)=8 6+14 =8 20=/=8 Dani is wrong